can you please help me with this problem? three times one number minus a second number is 8 and the sum of the number is 12. find the numbers.
Answer by stanbon(75887)

three times one number minus a second number is 8 and the sum of the number is 12. find the numbers.
------------------------------
Equations:
x + y = 12
3x-y = 8
----------------
Add the equations and solve for "x":
4x = 20
x = 5
-------------------
Substitute into x+y=12 and solve for "y":
5+y = 12
y = 7
================
Cheers,
Stan H.